OpenStudy (anonymous):

Which of the following correctly identifies the vertices that lie on the major axis of the conic section shown below? (x-2)^2/4+(y+5)^2/9=1 A. (-5,5) and (-5,-1) B. (0,-5) and (4,-5) C. (5,5) and (-1,-5) D. (2,-2) and (2,-8)

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

do you know what conic section this is?

OpenStudy (anonymous):

I think a hyperbola

OpenStudy (misty1212):

nope

OpenStudy (misty1212):

because it has a plus sign between the two terms, not a minus sign, it is not a hyperbola

OpenStudy (anonymous):

Ohh okay

OpenStudy (misty1212):

so what is it?

OpenStudy (anonymous):

ellipse

OpenStudy (misty1212):

right and what is the center of this ellipse?

OpenStudy (anonymous):

Im not sure but I think 4

OpenStudy (misty1212):

oh no the center is a point in the plane, an ordered pair \((h,k)\) not a number

OpenStudy (misty1212):

your ellipse is written in standard form \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] center is \((h,k)\)

OpenStudy (anonymous):

(-2,5)

OpenStudy (misty1212):

hmm close but no

OpenStudy (anonymous):

would it be (2,-5)

OpenStudy (misty1212):

\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\\ \frac{(x-2)^2}{4}+\frac{(y+5)^2}{9}=1\]

OpenStudy (misty1212):

yes \((2,-5)\)

OpenStudy (misty1212):

we need to know all these things to get the answer, i am not asking just to waste your time now we need to know two more things

OpenStudy (misty1212):



OpenStudy (anonymous):

the second one

OpenStudy (misty1212):

yes, and now we are just about done

OpenStudy (misty1212):

because we know the center is \((2,-5)\) and we know the major axis is vertical, so the first coordinates of both endpoints of the major axis will lie ABOVE and BELOW \((2,-5)\)

OpenStudy (misty1212):

that means this 

OpenStudy (misty1212):

we did not even have to find that they are 3 units above and below \((2,-5)\) because you only had one answer choice where the first coordinate was 2

OpenStudy (anonymous):

I think it would be A based off of that

OpenStudy (misty1212):

we could find the 3 if you like

OpenStudy (misty1212):

A???

OpenStudy (anonymous):

yes is that not right?

OpenStudy (misty1212):

i will let you look at the answer choices and the answer i wrote above and figure out which one it is for yourself